OpenStudy (anonymous):

Solve: 4(x + 1) = 4x + 10

OpenStudy (anonymous):

4x+4=4x+10 no answer

OpenStudy (anonymous):

thankz can you help me in another one if you dont mind ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

write girl

OpenStudy (anonymous):

this is algebra I

OpenStudy (anonymous):

Solve: 4(x − 1) = 2(x + 2) + 6